[Date Prev][Date Next][Thread Prev][Thread Next][Search] [Main Index] [Thread Index] [HEASARC Mailing List Archives]

CGRO biWeekly Status Report

  Compton Observatory Science Report #177, Friday February 17, 1995
      Chris Shrader, Compton Observatory Science Support Center

         Questions or comments can be sent to the CGRO SSC.
          Phone:    301/286-8434
          e-mail:   NSI_DECnet: GROSSC::SHRADER 
          Internet: shrader@grossc.gsfc.nasa.gov

                         Instrument Reports


EGRET operations were normal this biweekly period. Delivery of data
to the GRO SSC remains on schedule. Interaction with guest
investigators continues at a good level.

Work is nearing completion on the Second EGRET Catalog, and it should
be submitted to a journal shortly. Several abstracts have been
submitted for presentation at the American Physical Society Meeting
in April and at the International Cosmic Ray Conference. Careful
studies are in progress on the whole high energy gamma ray sky as
observed by EGRET including all the refinements that currently exist,
and they will be submitted as a series of papers over the next
several weeks.

At the moment, EGRET has just begun to view a region centered at
RA=98.5 and DEC=69.8, that includes QSO 0716+714.


OSSE detector #1 failed to step last night (16 Feb), the fourth such
event since launch and the first since 2 Jan. The motor drive
positioning process, which in normal operation moves the detectors
every 2 minutes, takes the detector off line when it detects a
positioning error. We restored the detector to normal operation
without any difficulty. How the drive knows to cause problems only in
the middle of the night remains a mystery.

OSSE is currently in normal operations on all four detectors. The
slewing response to BATSE burst triggers remains disabled while the
BATSE trigger is in its current low energy setting. We will re-enable
the slewing when the test period ends.

In viewing period 411.1 (14-21 Feb), the Z-axis target is Mrk 3
(Guest Investigator P. Nandra), and the X-axis target is Cas A (PI
team). Data from viewing periods 306 and 307 were delivered to the
Compton GRO Science Support Center archive this week. The targets
during these viewing periods were the Virgo region sky survey, the
galactic center region, and PSR 1800-21.


The COMPTEL instrument is performing well and continues routine

The collaboration forwarded to the SSC this week a major delivery of
archival data, encompassing all low-level and first high-level data
products for Phase 2 of the CGRO mission.

It has come to our attention that there is a page missing (page
number 8) from the COMPTEL preprint recently distributed on "COMPTEL
Observations of Galactic 26-Al Emission" by Diehl et al. A corrected
version of this preprint is currently being prepared for
redistribution. In the interim, postscript versions of the missing
page, or of the complete document, are available from either R. Diehl
at MPE (rod@mpe-garching.mpg.de) or via anonymous ftp to
unhgro.unh.edu (a user should "cd" to the /pub/papers directory).
(Note: This anonymous ftp facility is currently under construction;
it will eventually contain electronic postscript versions of all
COMPTEL team publications.)


The x-ray transient GRO J1719-24 = GRS 1716-249 remains visible at a
flux of 200-300 mCrab in the 20-100 keV energy range.  The LMXB
pulsar GX 1+4 has been markedly brighter in Earth occultation (about
300 mCrab 20-100 keV)  since it began to spin up.

The following sources were detected by the BATSE pulsed source
monitor in the past two weeks: Her X-1, Cen X-3, 4U 1626-67, OAO
1657-415, GX 1+4, Vela X-1, and GX 301-2. 

As of February 16, BATSE has detected 1228 cosmic gamma-ray bursts
out of a total of 3326 on-board triggers in 1395 days of operation.
There have been 748 triggers due to solar flares with emission above
60 keV.